OpenStudy (anonymous):

On Saturday, Louise ran 19/4 miles, Beth ran 45/10, Staci ran 4 3/5 miles. Which list shows these distances in order from greatest to least? a) 19/4 miles, 4 3/5 miles, 45/10 miles b) 4 3/5 miles, 45/10 miles, 19/4 miles c) 45/10 miles, 19/4 miles, 4 3/5 miles d) 45/10 miles, 4 3/5 miles, 19/4 miles

OpenStudy (mathmale):

I'd strongly suggest that you find the lowest common denominator of these several fractions. That would make it MUCH easier to arrange the factions in descending order.

OpenStudy (anonymous):

19/4 = 4.75 4 3/5 = 12/ 5 = 2.40 45/10 = 4.5 Just calculate the expression (divide or multiply then divide). This should help you get to the answer which is A.

OpenStudy (anonymous):

okay, thank you.